# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ 3.4 Definitions of Nonlinear Spectral Filters
Figure 1: Illustration of nonlinear spectral filters for equivariant machine learning on graphs. Given a graph $G$, the node features $\mathbf{X}$ are projected onto eigenspaces (analysis $\mathcal{A}$). The function $\Psi$ map a sequence of frequency responses to a sequence of frequency responses. The signal is represented in the graph domain using the nonlinear spectral coefficients (synthesis $\mathcal{S}$).
@@
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{ind})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (4)
$\Theta_{\text{val}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{val})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{K}\left[\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K}\right]_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (5)
where $\Psi:{\mathbb{R}}^{Jd}\rightarrow{\mathbb{R}}^{Jd}$ in $\Theta_{\text{ind}}$ and $\Psi:{\mathbb{R}}^{Kd}\rightarrow{\mathbb{R}}^{Kd}$ in $\Theta_{\text{val}}$. To adjust the feature output dimension at each node, we apply an MLP with shared weights to all nodes after the NLSF.
@@
$\Phi_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\text{\ \ and \ \ }\Phi_{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right).$ (6)
where $\Psi$ maps from ${\mathbb{R}}^{Jd}$ (resp. $\Phi_{\text{ind}}$) or ${\mathbb{R}}^{Kd}$ (resp. $\Phi_{\text{val}}$) to ${\mathbb{R}}^{d^{\prime}}$, and $d^{\prime}$ is the output dimension.
@@ 3.5 Laplacian Attention NLSFs
To understand which Laplacian and parameterization (index v/s value) of the NLSF are preferable in different settings, we follow the random geometric graph analysis outlined in [Paolino et al., 2023]. Specifically, we consider a setting where random geometric graphs are sampled from a metric-probability spaces $\mathcal{S}$. In such a case, the graph Laplacian approximates continuous Laplacians on the metric spaces under some conditions. We aim for our NLSFs to produce approximately the same outcome for any two graphs sampled from the same underlying metric space $\mathcal{S}$, ensuring that the NLSF is transferable. In App. A.5, we show that if the nodes of the graph are sampled uniformly from $\mathcal{S}$, then using the graph Laplacian $\mathbf{L}$ in Index NLSFs yields a transferable method. Conversely, if the nodes of the graph are sampled non-uniformly, and any two balls of the same radius in $\mathcal{S}$ have the same volume, then utilizing the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs is a transferable method.
Given that real-life graphs often fall between these two boundary cases, we present an architecture that chooses between the Index NLSFs and Value NLSFs, as illustrated in Fig. 2. Specifically, a soft attention mechanism is employed to dynamically choose between the two parameterizations, given by
@@
where $0\leq\alpha\leq 1$ is obtained using a softmax function to normalize the scores into attention weights, balancing each NLSFs’ contribution. It allows the model to adaptively select the most appropriate parameterization from the graph characteristics, promoting the transferability of the NLSFs.
@@ 4.1 Complexity of NLSFs
NLSFs are implemented by computing the eigenvectors of the GSO. Most existing spectral GNNs avoid direct eigendecomposition due to its perceived inefficiency. Instead, they use filters implemented by applying polynomials [Kipf and Welling, 2016, Defferrard et al., 2016a] or rational functions [Levie et al., 2018, Bianchi et al., 2021] to the GSO in the spatial domain. However, power iteration-based eigendecomposition algorithms, e.g., variants of the Lanczos method, can be highly efficient [Saad, 2011, Lanczos, 1950]. For matrices with $E$ non-zero entries, the computational complexity of one iteration for finding $J$ eigenvectors corresponding to the smallest or largest eigenvalues (called leading eigenvectors) is $O(JE)$. In practice, these eigendecomposition algorithms converge quickly due to their super-exponential convergence rate, often requiring only a few iterations.
This makes NLSFs applicable to node-level tasks on large sparse graphs, as they rely solely on the leading eigenvectors. In Sec. 4.4, we show that using the leading eigenvectors can approximate GSO well in the context of learning on graphs. For node-level tasks, such as semi-supervised node classification, the leading eigenvectors only need to be pre-computed once, with a complexity of $O(JE)$. During the learning phase, each step of the architecture search and hyperparameter optimization takes $O(NJd)$ complexity for analysis and synthesis, and $O(J^{2}d^{2})$ for the MLP in the spectral domain, which is faster than the complexity $O(Ed^{2})$ of message passing or standard spectral methods if $NJ<Ed$. Detailed empirical studies on runtime analysis are provided in the App. C.
@@ 4.2 Equivariance of Node-level NLSFs
We demonstrate the node-level equivariance of our NLSFs, ensuring that our method respects the functional shift symmetries. The proof is given in App. A.1.
@@
Index NLSFs in Eq. (4) are equivariant to the graph functional shift $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@ 4.3 Universal Approximation and Expressivity
Node-Level Universal Approximation. We begin with a setting where a graph is given as a fixed domain, and the data distribution consists of multiple signals defined on this graph. An example of this setup is a spatiotemporal graph [Cini and Marisca, 2022], e.g., traffic networks, where a fixed sensor system defines a graph and the different signals represent the sensor readings collected at different times.
In App. A.2.1, we prove the following lemma, which shows that linear NLSFs exhaust the space of linear operators that commute with graph functional shifts.
@@
Suppose that ${\scriptstyle\sum}_{j}\mathbf{P}_{j}={\scriptstyle\sum}_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$. A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commute with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
Lemma 4.1 shows a close relationship between functions that commute with functional shifts and those defined in the spectral domain. This motivates the following construction of a pseudo-metric on ${\mathbb{R}}^{N\times d}$. In the case of relaxed functional shifts, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a signal pseudo-metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined by
@@
This pseudo metric can be made into a metric by considering each equivalence class of signals with zero distance as a single point in the space. As MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{K\times d}$ (the universal approximation theorem [Cybenko, 1989, Funahashi, 1989, Leshno et al., 1993]), node-level NLSFs can approximate any continuous function that maps (equivalence classes of) signals to (equivalence classes of) signals. For details, see App. A.2.2. A similar analysis applies to hard functional shifts.
Graph-Level Universal Approximation. The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we consider again the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We define the distance between any two graphs with GSOs and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ to be
@@
This definition can be extended into a metric by considering the space $\mathcal{G}$ of equivalence classes of graph-signals with distance 0. As before, this distance inherits the universal approximation properties of standard MLPs. Namely, any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by NLSFs based on MLPs. Additional details are in App. A.2.3.
Graph-Level Expressivity of Pooling-NLSFs. In App. A.2.4, we show that Pooling-NLSFs are more expressive than graph-level NLSF when any $L_{p}$ norm is used in Eq. (4) and Eq. (5) with $p\neq 2$, both in the definition of the NLSF and as the pooling method. Specifically, for every graph-level NLSF, there is a Pooling-NLSF that coincides with it. Additionally, there are graph signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ for which a Pooling-NLSF can attain different values, whereas any graph-level NLSF must attain the same value. Hence, Pooling-NLSFs have improved discriminative power compared to graph-level NLSFs. Indeed, as shown in Tab. LABEL:tab:graph_classification_accuracy_vertical, Pooling-NLSFs outperform Graph-level NLSFs in practice, which can be attributed to their increased expressivity.
@@ 4.4 Uniform Approximation of GSOs by Their Leading Eigenvectors
Since NLSFs on large graphs are based on the leading eigenvectors of ${\bm{\Delta}}$, we justify its low-rank approximation in the following. While approximating matrices with low-rank matrices might lead to a high error in the spectral and Frobenius norms, we show that such an approximation entails a uniformly small error in the cut norm. We define and interpret the cut norm in App. A.4.1, and explain why it is a natural graph similarity measure for graph machine learning.
The following theorem is a corollary of the Constructive Weak Regularity Lemma presented in [Finkelshtein et al., 2024]. Its proof is presented in App. A.4.
@@
Note that the bound in Thm. 4.1 is uniformly small, independently of $\mathbf{M}$ and its dimension $N$. This theorem justifies using the leading eigenvectors when working with the adjacency matrix as the GSO. For a justification when working with other GSOs see App. A.4.
@@ 5 Experimental Results
We evaluate the NLSFs on node and graph classification tasks. Additional implementation details are in App. B. Additional experiments, including runtime analysis and ablation studies, are in App. C.
@@ 5.1 Semi-Supervised Node Classification
$\mathtt{att}$-Node-level NLSFs 84.75$\pm$0.7 73.62$\pm$1.1 81.93$\pm$1.0 49.68$\pm$1.6 38.25$\pm$0.7 34.72$\pm$0.9
We first demonstrate the main advantage of the proposed Node-level NLSFs over existing GNNs with convolution design on semi-supervised node classification tasks. We test three citation networks [Sen et al., 2008, Yang et al., 2016]: Cora, Citeseer, and Pubmed. In addition, we explore three heterophilic graphs: Chameleon, Squirrel, and Actor [Rozemberczki et al., 2021, Tang et al., 2009]. For more comprehensive descriptions of these datasets, see App. B.
We compare the Node-level NLSFs using Laplacian attention with existing spectral GNNs for node-level predictions, including GCN [Kipf and Welling, 2016], ChebNet [Defferrard et al., 2016a], ChebNetII [He et al., 2022], CayleyNet [Levie et al., 2018], APPNP [Klicpera et al., 2019], GPRGNN [Chien et al., 2020], and ARMA [Bianchi et al., 2021]. We also consider GAT [Veličković et al., 2017] and SAGE [Hamilton et al., 2017]. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we apply the standard splits following [Yang et al., 2016], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we use the sparse splitting as in [Chien et al., 2020, He et al., 2022], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. We measure the classification quality by computing the average classification accuracy with a 95% confidence interval over 10 random splits. We utilize the source code released by the authors for the baseline algorithms and optimize their hyperparameters using Optuna [Akiba et al., 2019]. Each model’s hyperparameters are fine-tuned to achieve the highest possible accuracy. Detailed hyperparameter settings are provided in App. B.
Tab. LABEL:tab:node_classification_accuracy presents the node classification accuracy of our Node-level NLSFs using Laplacian attention and the various competing baselines. We see that $\mathtt{att}$-Node-level NLSFs outperform the competing models on the Cora, Citeseer, and Chameleon datasets. Notably, it shows remarkable performance on the densely connected Squirrel graph, outperforming the baselines by a large margin. This can be explained by the sparse version in Eq. (9) of Thm. 4.1, which shows that the denser the graphs is, the better its rank-$J$ approximation. For the Pubmed and Actor datasets, $\mathtt{att}$-Node-level NLSFs yield the second-best results, which are comparable to the best results obtained by APPNP.

### v2

@@ 3.4 Definitions of Nonlinear Spectral Filters
Figure 1: Illustration of nonlinear spectral filters for equivariant machine learning on graphs. Given a graph $G$, the node features $\mathbf{X}$ are projected onto eigenspaces (analysis $\mathcal{A}$). The function $\Psi$ map a sequence of frequency coefficients to a sequence of frequency coefficients. The coefficients are synthesized to the graph domain using the using $\mathcal{S}$.
@@
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{ind})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi_{\text{ind}}(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))=\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J+1},$ (4)
$\Theta_{\text{val}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{val})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi_{\text{val}}(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))=\left[\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{K+1}\left[\Psi_{\text{val}}\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{K+1},$ (5)
where $\Psi_{\text{ind}}:{\mathbb{R}}^{(J+1)d}\rightarrow{\mathbb{R}}^{(J+1)d\times(J+1)\widetilde{d}}$ and $\Psi_{\text{val}}:{\mathbb{R}}^{(K+1)d}\rightarrow{\mathbb{R}}^{(K+1)d\times(K+1)\widetilde{d}}$ are called nonlinear frequency responses, and $\widetilde{d}$ is the output dimension. To adjust the feature output dimension, we apply an MLP with shared weights to all nodes after the NLSF. In the case when $\widetilde{d}=d$ and the filters operator diagonally (i.e., the product and division are element-wise in synthesis), we refer to it as diag-NLSF. See App. C for more details.
@@
$\Phi_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\widehat{\Psi}_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\text{\ \ and \ \ }\Phi_{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\widehat{\Psi}_{\text{val}}\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right),$ (6)
where $\widehat{\Psi}_{\text{ind}}:{\mathbb{R}}^{(J+1)d}\rightarrow{\mathbb{R}}^{d^{\prime}}$, $\widehat{\Psi}_{\text{val}}:{\mathbb{R}}^{(K+1)d}\rightarrow{\mathbb{R}}^{d^{\prime}}$, and $d^{\prime}$ is the output dimension.
@@ 3.5 Laplacian Attention NLSFs
To understand which Laplacian and parameterization (index v/s value) of the NLSF are preferable in different settings, we follow the random geometric graph analysis outlined in [74]. Specifically, we consider a setting where random geometric graphs are sampled from a metric-probability space $\mathcal{S}$. In such a case, the graph Laplacian approximates continuous Laplacians on the metric spaces under some conditions. We aim for our NLSFs to produce approximately the same outcome for any two graphs sampled from the same underlying metric space $\mathcal{S}$, ensuring that the NLSF is transferable. In App. D.5, we show that if the nodes of the graph are sampled uniformly from $\mathcal{S}$, then using the graph Laplacian $\mathbf{L}$ in Index NLSFs yields a transferable method. Conversely, if the nodes of the graph are sampled non-uniformly, and any two balls of the same radius in $\mathcal{S}$ have the same volume, then utilizing the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs is a transferable method. Given that graphs may fall between these two boundary cases, we present an architecture that chooses between the Index NLSFs with respect to $\mathbf{L}$ and Value NLSFs with respect to $\mathbf{N}$, as illustrated in Fig. 2. While the above theoretical setting may not be appropriate as a model for every real-life graph dataset, it suggests that index NLSF may be more appropriate with $\mathbf{L}$, value NLSFs with $\mathbf{N}$, and different graphs are more appropriately analyzed by different balances between these two cases.
In the Laplacian attention architecture, a soft attention mechanism is employed to dynamically choose between the two parameterizations, given by
@@
where $0\leq\alpha\leq 1$ is obtained using a softmax function to normalize the scores into attention weights, balancing each NLSFs’ contribution.
@@ 4.1 Complexity of NLSFs
NLSFs are implemented by computing the eigenvectors of the GSO. Most existing spectral GNNs avoid direct eigendecomposition due to its perceived inefficiency. Instead, they use filters implemented by applying polynomials [49, 22] or rational functions [61, 5] to the GSO in the spatial domain. However, power iteration-based eigendecomposition algorithms, e.g., variants of the Lanczos method, can be highly efficient [80, 55]. For matrices with $E$ non-zero entries, the computational complexity of one iteration for finding $J$ eigenvectors corresponding to the smallest or largest eigenvalues (called leading eigenvectors) is $O(JE)$. In practice, these eigendecomposition algorithms converge quickly due to their super-exponential convergence rate, often requiring only a few iterations, which makes them as efficient as message passing networks of signals with $\sqrt{J}$ channels.
This makes NLSFs applicable to node-level tasks on large sparse graphs, as demonstrated empirically in App. F.5, since they rely solely on the leading eigenvectors. In Sec. 4.4, we show that using the leading eigenvectors can approximate GSOs well in the context of learning on graphs. Note that we can precompute the spectral projections of the signal before training. For node-level tasks, such as semi-supervised node classification, the leading eigenvectors only need to be pre-computed once, with a complexity of $O(JE)$. This During the learning phase, each step of the architecture search and hyperparameter optimization takes $O(NJd)$ complexity for analysis and synthesis, and $O(J^{2}d^{2})$ for the MLP in the spectral domain, which is faster than the complexity $O(Ed^{2})$ of message passing or standard spectral methods if $NJ<Ed$. Empirical studies on runtime analysis are in App. F.
@@ 4.2 Equivariance of Node-level NLSFs
We demonstrate the node-level equivariance of our NLSFs, ensuring that our method respects the functional shift symmetries. The proof is given in App. D.1.
@@
Index NLSFs in Eq. (4) are equivariant to the graph functional shifts $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@ 4.3 Universal Approximation and Expressivity
Node-Level Universal Approximation. We begin with a setting where a graph is given as a fixed domain, and the data distribution consists of multiple signals defined on this graph. An example of this setup is a spatiotemporal graph [17], e.g., traffic networks, where a fixed sensor system defines a graph and the different signals represent the sensor readings collected at different times.
In App. D.2.1, we prove the following lemma, which shows that linear NLSFs exhaust the space of linear operators that commute with graph functional shifts.
@@
A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commutes with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
Lemma 4.1 shows a close relationship between functions that commute with functional shifts and those defined in the spectral domain. This motivates the following construction of a pseudo-metric on ${\mathbb{R}}^{N\times d}$. In the case of relaxed functional shifts, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{(K+1)\times d}$. We pull back the Euclidean metric to the spatial domain to define a signal pseudo-metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined by
@@
This pseudo metric can be made into a metric by considering each equivalence class of signals with zero distance as a single point in the space. As MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{(K+1)\times d}\rightarrow{\mathbb{R}}^{(K+1)\times d}$ (the universal approximation theorem [19, 34, 58]), node-level NLSFs can approximate any continuous function that maps (equivalence classes of) signals to (equivalence classes of) signals. For details, see App. D.2.2. A similar analysis applies to hard functional shifts.
Graph-Level Universal Approximation. The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we consider again the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{(K+1)\times d}$. We define the distance between any two graphs with GSOs and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ to be
@@
This definition can be extended into a metric by considering the space $\mathcal{G}$ of equivalence classes of graph-signals with distance 0. As before, this distance inherits the universal approximation properties of standard MLPs. Namely, any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by NLSFs based on MLPs. Additional details are in App. D.2.3.
Graph-Level Expressivity of Pooling-NLSFs. In App. D.2.4, we show that Pooling-NLSFs are more expressive than graph-level NLSF when any $L_{p}$ norm is used in Eq. (4) and Eq. (5) with $p\neq 2$, both in the definition of the NLSF and as the pooling method. Specifically, for every graph-level NLSF, there is a Pooling-NLSF that coincides with it. Additionally, there are graph signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ for which a Pooling-NLSF can attain different values, whereas any graph-level NLSF must attain the same value. Hence, Pooling-NLSFs have improved discriminative power compared to graph-level NLSFs. Indeed, as shown in Tab. LABEL:tab:graph_classification_accuracy_vertical, Pooling-NLSFs outperform Graph-level NLSFs in practice, which can be attributed to their increased expressivity. We refer to App. D.2.5 for additional discussion on graph-level expressivity.
@@ 4.4 Uniform Approximation of GSOs by Their Leading Eigenvectors
Since NLSFs on large graphs are based on the leading eigenvectors of ${\bm{\Delta}}$, we justify its low-rank approximation in the following. While approximating matrices with low-rank matrices might lead to a high error in the spectral and Frobenius norms, we show that such an approximation entails a uniformly small error in the cut norm. We define and interpret the cut norm in App. D.4.1, and explain why it is a natural graph similarity measure for graph machine learning.
The following theorem is a corollary of the Constructive Weak Regularity Lemma presented in [29]. Its proof is presented in App. D.4.
@@
Note that the bound in Thm. 4.1 is uniformly small, independently of $\mathbf{M}$ and its dimension $N$. This theorem justifies using the leading eigenvectors when working with the adjacency matrix as the GSO. For a justification when working with other GSOs see App. D.4.
@@ 5 Experimental Results
We evaluate the NLSFs on node and graph classification tasks. Additional implementation details are in App. E, and additional experiments, including runtime analysis and ablation studies, are in App. F.
@@ 5.1 Semi-Supervised Node Classification
JacobiConv 84.12$\pm$0.7 72.59$\pm$1.4 82.05$\pm$1.9 49.66$\pm$1.9 33.65$\pm$0.8 34.61$\pm$0.7
BernNet 82.96$\pm$1.1 71.25$\pm$1.0 81.07$\pm$1.6 42.65$\pm$3.4 31.68$\pm$1.5 33.92$\pm$0.8
Specformer 82.27$\pm$0.7 73.45$\pm$1.4 81.62$\pm$1.0 49.79$\pm$1.2 38.24$\pm$0.9 34.12$\pm$0.6
OptBasisGNN 81.97$\pm$1.2 70.46$\pm$1.6 80.38$\pm$0.9 47.12$\pm$2.4 37.66$\pm$1.1 34.84$\pm$1.3
$\mathtt{att}$-Node-level NLSFs 85.37$\pm$1.8 75.41$\pm$0.8 82.22$\pm$1.2 50.58$\pm$1.3 38.39$\pm$0.9 35.13$\pm$1.0
We first demonstrate the main advantage of the proposed Node-level NLSFs over existing GNNs with convolution design on semi-supervised node classification tasks. We test three citation networks [84, 100]: Cora, Citeseer, and Pubmed. In addition, we explore three heterophilic graphs: Chameleon, Squirrel, and Actor [79, 90]. For more comprehensive descriptions of these datasets, see App. E.
We compare the Node-level NLSFs using Laplacian attention with existing spectral GNNs for node-level predictions, including GCN [49], ChebNet [22], ChebNetII [41], CayleyNet [61], APPNP [50], GPRGNN [16], ARMA [5], JacobiConv [96], BernNet [42], Specformer [7], and OptBasisGNN [38]. We also consider GAT [93] and SAGE [39]. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we apply the standard splits following [100], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we use the sparse splitting as in [16, 41], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. We measure the classification quality by computing the average classification accuracy with a 95% confidence interval over 10 random splits. We utilize the source code released by the authors for the baseline algorithms and optimize their hyperparameters using Optuna [1]. Each model’s hyperparameters are fine-tuned to achieve the highest possible accuracy. Detailed hyperparameter settings are provided in App. E.
Tab. LABEL:tab:node_classification_accuracy presents the node classification accuracy of our NLSFs using Laplacian attention and the various competing baselines. We see that $\mathtt{att}$-Node-level NLSFs outperform the competing models on the Cora, Citeseer, and Chameleon datasets. Notably, it shows remarkable performance on the densely connected Squirrel graph, outperforming the baselines by a large margin. This can be explained by the sparse version in Eq. (21) of Thm. 4.1, which shows that the denser the graphs is, the better its rank-$J$ approximation. For the Pubmed and Actor datasets, $\mathtt{att}$-Node-level NLSFs yield the second-best results, which are comparable to the best results obtained by APPNP.